OpenStudy (anonymous):

2q/(2q+3) - 2q/(2q-3) = 1

OpenStudy (anonymous):

solve for q?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oopes, messed all that up right quick

OpenStudy (anonymous):

\[\frac{2q ( 2q - 3)- 2q (2q+3)}{(2q-3)(2q+3)} =1\] \[\frac {2q ((2q - 3) - (2q +3))}{ 4q^{2} - 9}=1\]

OpenStudy (anonymous):

does that help at all / do you see where that comes from?

OpenStudy (anonymous):

yes I understand so far

OpenStudy (anonymous):

Can you simplify that further?

OpenStudy (anonymous):

I believe so. thank you for your help!

OpenStudy (anonymous):

:) welcome